OpenStudy (anonymous):

solve the system by using elimination or elimination with multiplication. x-y=3 x+y=7

OpenStudy (anonymous):

x - y = 3 and x + y = 7 add these together to get, 2x=10, so x=5 plug this value into one of the 1st equations 5+y=7 y=7-5, y=2 (x,y)=(5, 2)

OpenStudy (anonymous):

x =5 and y=2